\begin{thebibliography}{10}

\bibitem{loneliness}
Social isolation and loneliness.
\newblock \url{https://www.who.int/teams/social-determinants-of-health/demographic-change-and-healthy-ageing/social-isolation-and-loneliness}.
\newblock Accessed: 2025-08.

\bibitem{DatgControl}
Data controls faq.
\newblock \url{https://help.openai.com/en/articles/7730893-data-controls-faq}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/OOaHO}.

\bibitem{leaderboard}
Leaderboard overview.
\newblock \url{https://huggingface.co/spaces/lmarena-ai/lmarena-leaderboard}, 2025.
\newblock Accessed: 2025-08.

\bibitem{anderljung2024protecting}
Markus Anderljung, Julian Hazell, and Moritz von Knebel.
\newblock Protecting society from ai misuse: when are restrictions on capabilities warranted?
\newblock {\em AI \& SOCIETY}, pages 1--17, 2024.

\bibitem{anderson2001taxonomy}
Lorin~W Anderson and David~R Krathwohl.
\newblock {\em A taxonomy for learning, teaching, and assessing: A revision of Bloom's taxonomy of educational objectives: complete edition}.
\newblock Addison Wesley Longman, Inc., 2001.

\bibitem{bastani2024generative}
Hamsa Bastani, Osbert Bastani, Alp Sungu, Haosen Ge, Ozge Kabakc{\i}, and Rei Mariman.
\newblock Generative ai can harm learning.
\newblock {\em Available at SSRN}, 4895486, 2024.

\bibitem{Jobs}
Nick Bilton.
\newblock Steve jobs was a low-tech parent.
\newblock \url{https://www.nytimes.com/2014/09/11/fashion/steve-jobs-apple-was-a-low-tech-parent.html}, 2014.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/61YCq}.

\bibitem{dataCenter}
Eliot Brown.
\newblock Openai commits to giant u.a.e. data center in global expansion.
\newblock \url{https://www.wsj.com/tech/open-ai-abu-dhabi-data-center-1c3e384d}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/218rI}.

\bibitem{CasesDatabase}
Damien Charlotin.
\newblock Ai hallucination cases.
\newblock \url{https://www.damiencharlotin.com/hallucinations/}, 2025.
\newblock Accessed: 2025-07.

\bibitem{energy}
Sophia Chen.
\newblock How much energy will ai really consume? the good, the bad and the unknown.
\newblock \url{https://www.nature.com/articles/d41586-025-00616-z}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/pEthw}.

\bibitem{chuang2023dola}
Yung-Sung Chuang, Yujia Xie, Hongyin Luo, Yoon Kim, James Glass, and Pengcheng He.
\newblock Dola: Decoding by contrasting layers improves factuality in large language models.
\newblock {\em arXiv preprint arXiv:2309.03883}, 2023.

\bibitem{das2025security}
Badhan~Chandra Das, M~Hadi Amini, and Yanzhao Wu.
\newblock Security and privacy challenges of large language models: A survey.
\newblock {\em ACM Computing Surveys}, 57(6):1--39, 2025.

\bibitem{de2023chatgpt}
Luigi De~Angelis, Francesco Baglivo, Guglielmo Arzilli, Gaetano~Pierpaolo Privitera, Paolo Ferragina, Alberto~Eugenio Tozzi, and Caterina Rizzo.
\newblock Chatgpt and the rise of large language models: the new ai-driven infodemic threat in public health.
\newblock {\em Frontiers in public health}, 11:1166120, 2023.

\bibitem{dergaa2024tools}
Ismail Dergaa, Helmi Ben~Saad, Jordan~M Glenn, Badii Amamou, Mohamed Ben~Aissa, Noomen Guelmami, Feten Fekih-Romdhane, and Karim Chamari.
\newblock From tools to threats: a reflection on the impact of artificial-intelligence chatbots on cognitive health.
\newblock {\em Frontiers in psychology}, 15:1259845, 2024.

\bibitem{eloundou2023gpts}
Tyna Eloundou, Sam Manning, Pamela Mishkin, and Daniel Rock.
\newblock Gpts are gpts: An early look at the labor market impact potential of large language models.
\newblock {\em arXiv preprint arXiv:2303.10130}, 2023.

\bibitem{gmail_train_AI}
Geoffrey~A. Fowler.
\newblock Your instagrams are training ai. there’s little you can do about it.
\newblock \url{https://www.washingtonpost.com/technology/2023/09/08/gmail-instagram-facebook-trains-ai/}, 2023.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/4kJTD}.

\bibitem{metaPersonalize}
Geoffrey~A. Fowler.
\newblock Zuckerberg’s new meta ai app gets personal in a very creepy way.
\newblock \url{https://www.washingtonpost.com/technology/2025/05/05/meta-ai-privacy/}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/8VQd9}.

\bibitem{gallegos2024bias}
Isabel~O Gallegos, Ryan~A Rossi, Joe Barrow, Md~Mehrab Tanjim, Sungchul Kim, Franck Dernoncourt, Tong Yu, Ruiyi Zhang, and Nesreen~K Ahmed.
\newblock Bias and fairness in large language models: A survey.
\newblock {\em Computational Linguistics}, 50(3):1097--1179, 2024.

\bibitem{teensAIFriend}
JOCELYN GECKER.
\newblock Teens say they are turning to ai for friendship.
\newblock \url{https://apnews.com/article/ai-companion-generative-teens-mental-health-9ce59a2b250f3bd0187a717ffa2ad21f}, 2025.
\newblock Accessed: 2025-08.

\bibitem{gekhman2024does}
Zorik Gekhman, Gal Yona, Roee Aharoni, Matan Eyal, Amir Feder, Roi Reichart, and Jonathan Herzig.
\newblock Does fine-tuning llms on new knowledge encourage hallucinations?
\newblock {\em arXiv preprint arXiv:2405.05904}, 2024.

\bibitem{gerlich2025ai}
Michael Gerlich.
\newblock Ai tools in society: Impacts on cognitive offloading and the future of critical thinking.
\newblock {\em Societies}, 15(1):6, 2025.

\bibitem{MAHA}
Caitlin Gilbert, Emily Wright, Fenit Nirappil, and Lauren Weber.
\newblock The maha report’s ai fingerprints, annotated.
\newblock \url{https://www.washingtonpost.com/health/2025/05/30/maha-report-ai-white-house/}, 2025.
\newblock Accessed: 2025-07. Archived at \url{https://archive.ph/cPUpY}.

\bibitem{gonzalez2025reinforcement}
Kristian Gonz{\'a}lez~Barman, Simon Lohse, and Henk~W de~Regt.
\newblock Reinforcement learning from human feedback in llms: Whose culture, whose values, whose perspectives?
\newblock {\em Philosophy \& Technology}, 38(2):1--26, 2025.

\bibitem{herbener2025lonely}
Arthur~Bran Herbener and Malene~Flensborg Damholdt.
\newblock Are lonely youngsters turning to chatbots for companionship? the relationship between chatbot usage and social connectedness in danish high-school students.
\newblock {\em International Journal of Human-Computer Studies}, 196:103409, 2025.

\bibitem{delusion_nyt}
Kashmir Hill.
\newblock They asked an a.i. chatbot questions. the answers sent them spiraling.
\newblock \url{https://www.nytimes.com/2025/06/13/technology/chatgpt-ai-chatbots-conspiracies.html}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/5MPaG}.

\bibitem{hochreiter1997long}
Sepp Hochreiter and J{\"u}rgen Schmidhuber.
\newblock Long short-term memory.
\newblock {\em Neural computation}, 9(8):1735--1780, 1997.

\bibitem{hofmann2024ai}
Valentin Hofmann, Pratyusha~Ria Kalluri, Dan Jurafsky, and Sharese King.
\newblock Ai generates covertly racist decisions about people based on their dialect.
\newblock {\em Nature}, 633(8028):147--154, 2024.

\bibitem{huang2025survey}
Lei Huang, Weijiang Yu, Weitao Ma, Weihong Zhong, Zhangyin Feng, Haotian Wang, Qianglong Chen, Weihua Peng, Xiaocheng Feng, Bing Qin, et~al.
\newblock A survey on hallucination in large language models: Principles, taxonomy, challenges, and open questions.
\newblock {\em ACM Transactions on Information Systems}, 43(2):1--55, 2025.

\bibitem{delusion_wsj}
Julie Jargon.
\newblock He had dangerous delusions. chatgpt admitted it made them worse.
\newblock \url{https://www.wsj.com/tech/ai/chatgpt-chatbot-psychology-manic-episodes-57452d14}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/gsavP}.

\bibitem{jindal2022misguided}
Atin Jindal.
\newblock Misguided artificial intelligence: how racial bias is built into clinical models.
\newblock {\em The Brown journal of hospital medicine}, 2(1):38021, 2022.

\bibitem{jurafsky2000speech}
Dan Jurafsky.
\newblock {\em Speech \& language processing}.
\newblock Pearson Education India, 2000.

\bibitem{kaplan2020scaling}
Jared Kaplan, Sam McCandlish, Tom Henighan, Tom~B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei.
\newblock Scaling laws for neural language models.
\newblock {\em arXiv preprint arXiv:2001.08361}, 2020.

\bibitem{kaufmann2024survey}
Timo Kaufmann, Paul Weng, Viktor Bengs, and Eyke H{\"u}llermeier.
\newblock A survey of reinforcement learning from human feedback.
\newblock 2024.

\bibitem{kim2025medical}
Yubin Kim, Hyewon Jeong, Shan Chen, Shuyue~Stella Li, Mingyu Lu, Kumail Alhamoud, Jimin Mun, Cristina Grau, Minseok Jung, Rodrigo Gameiro, et~al.
\newblock Medical hallucinations in foundation models and their impact on healthcare.
\newblock {\em arXiv preprint arXiv:2503.05777}, 2025.

\bibitem{kojima2022large}
Takeshi Kojima, Shixiang~Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa.
\newblock Large language models are zero-shot reasoners.
\newblock {\em Advances in neural information processing systems}, 35:22199--22213, 2022.

\bibitem{kosmyna2025your}
Nataliya Kosmyna, Eugene Hauptmann, Ye~Tong Yuan, Jessica Situ, Xian-Hao Liao, Ashly~Vivian Beresnitzky, Iris Braunstein, and Pattie Maes.
\newblock Your brain on chatgpt: Accumulation of cognitive debt when using an ai assistant for essay writing task.
\newblock {\em arXiv preprint arXiv:2506.08872}, 2025.

\bibitem{kotek2023gender}
Hadas Kotek, Rikker Dockum, and David Sun.
\newblock Gender bias and stereotypes in large language models.
\newblock In {\em Proceedings of the ACM collective intelligence conference}, pages 12--24, 2023.

\bibitem{lambert2025reinforcement}
Nathan Lambert.
\newblock Reinforcement learning from human feedback.
\newblock {\em arXiv preprint arXiv:2504.12501}, 2025.

\bibitem{chatgpt_leak}
Frank Landymore.
\newblock Leaked chatgpt conversation shows user identified as lawyer asking how to ``displace a small amazonian indigenous community from their territories in order to build a dam and a hydroelectric plant".
\newblock \url{https://futurism.com/leaked-chatgpt-lawyer-displace-amazonian}, 2025.
\newblock Accessed: 2025-08.

\bibitem{lee2025impact}
Hao-Ping Lee, Advait Sarkar, Lev Tankelevitch, Ian Drosos, Sean Rintel, Richard Banks, and Nicholas Wilson.
\newblock The impact of generative ai on critical thinking: Self-reported reductions in cognitive effort and confidence effects from a survey of knowledge workers.
\newblock In {\em Proceedings of the 2025 CHI conference on human factors in computing systems}, pages 1--22, 2025.

\bibitem{liu2025illusion}
Tingting Liu, Salvatore Giorgi, Ankit Aich, Allison Lahnala, Brenda Curtis, Lyle Ungar, and Jo{\~a}o Sedoc.
\newblock The illusion of empathy: How ai chatbots shape conversation perception.
\newblock In {\em Proceedings of the AAAI Conference on Artificial Intelligence}, volume~39, pages 14327--14335, 2025.

\bibitem{macnamara2024does}
Brooke~N Macnamara, Ibrahim Berber, M~Cenk {\c{C}}avu{\c{s}}o{\u{g}}lu, Elizabeth~A Krupinski, Naren Nallapareddy, Noelle~E Nelson, Philip~J Smith, Amy~L Wilson-Delfosse, and Soumya Ray.
\newblock Does using artificial intelligence assistance accelerate skill decay and hinder skill development without performers’ awareness?
\newblock {\em Cognitive Research: Principles and Implications}, 9(1):46, 2024.

\bibitem{pewResearch}
Colleen McClain, Brian Kennedy, Jeffrey Gottfried, Monica Anderson, and Giancarlo Pasquini.
\newblock How the u.s. public and ai experts view artificial intelligence.
\newblock \url{https://www.pewresearch.org/internet/2025/04/03/how-the-us-public-and-ai-experts-view-artificial-intelligence/}, 2025.
\newblock Accessed: 2025-08.

\bibitem{mihalcea2025ai}
Rada Mihalcea, Oana Ignat, Longju Bai, Angana Borah, Luis Chiruzzo, Zhijing Jin, Claude Kwizera, Joan Nwatu, Soujanya Poria, and Thamar Solorio.
\newblock Why ai is weird and shouldn't be this way: Towards ai for everyone, with everyone, by everyone.
\newblock In {\em Proceedings of the AAAI Conference on Artificial Intelligence}, volume~39, pages 28657--28670, 2025.

\bibitem{isItSafe}
Heidi Mitchell.
\newblock Is it safe to share personal information with a chatbot?
\newblock \url{https://www.wsj.com/tech/ai/ai-chatbot-sharing-personal-information-229d41a0}, 2024.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/Okgix}.

\bibitem{sycophancy}
Heidi Mitchell.
\newblock That chatbot may just be telling you what you want to hear.
\newblock \url{https://www.wsj.com/tech/ai/ai-chatbot-agree-flatter-users-1787e1a7}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/zMcXY}.

\bibitem{moore2025expressing}
Jared Moore, Declan Grabb, William Agnew, Kevin Klyman, Stevie Chancellor, Desmond~C Ong, and Nick Haber.
\newblock Expressing stigma and inappropriate responses prevents llms from safely replacing mental health providers.
\newblock In {\em Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency}, pages 599--627, 2025.

\bibitem{motoki2024more}
Fabio Motoki, Valdemar Pinho~Neto, and Victor Rodrigues.
\newblock More human than human: measuring chatgpt political bias.
\newblock {\em Public Choice}, 198(1):3--23, 2024.

\bibitem{motoki2025assessing}
Fabio~YS Motoki, Valdemar~Pinho Neto, and Victor Rangel.
\newblock Assessing political bias and value misalignment in generative artificial intelligence.
\newblock {\em Journal of Economic Behavior \& Organization}, page 106904, 2025.

\bibitem{deepseek_leak}
Gal Nagli.
\newblock Wiz research uncovers exposed deepseek database leaking sensitive information, including chat history.
\newblock \url{https://www.wiz.io/blog/wiz-research-uncovers-exposed-deepseek-database-leak}, 2025.
\newblock Accessed: 2025-08.

\bibitem{AI_water}
Leonardo Nicoletti, Michelle Ma, and Dina Bass.
\newblock Ai is draining water from areas that need it most.
\newblock \url{https://www.bloomberg.com/graphics/2025-ai-impacts-data-centers-water-data/}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/QXNJi}.

\bibitem{job_wsj}
Peggy Noonan.
\newblock Ai is here, and a quiet havoc has begun.
\newblock \url{https://www.wsj.com/opinion/ai-is-here-and-a-quiet-havoc-has-begun-15322dd7}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/E3kxj}.

\bibitem{companionAddiction}
James O'Donnell.
\newblock Ai companions are the final stage of digital addiction, and lawmakers are taking aim.
\newblock \url{https://www.technologyreview.com/2025/04/08/1114369/ai-companions-are-the-final-stage-of-digital-addiction-and-lawmakers-are-taking-aim/}, 2025.
\newblock Accessed: 2025-08.

\bibitem{woke_AI}
Will Oremus.
\newblock Trump is targeting ‘woke ai.’ here’s what that means.
\newblock \url{https://www.washingtonpost.com/technology/2025/07/24/trump-ai-woke-executive-order/}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/yiTL1}.

\bibitem{cognitiveCost}
David Ovalle.
\newblock Is ai rewiring our minds? scientists probe cognitive cost of chatbots.
\newblock \url{https://www.washingtonpost.com/health/2025/06/29/chatgpt-ai-brain-impact/}, 2025.
\newblock Accessed: 2025-07. Archived at \url{https://archive.ph/xNwWZ}.

\bibitem{park2024ai}
Peter~S Park, Simon Goldstein, Aidan O’Gara, Michael Chen, and Dan Hendrycks.
\newblock Ai deception: A survey of examples, risks, and potential solutions.
\newblock {\em Patterns}, 5(5), 2024.

\bibitem{Gates}
Emily Retter.
\newblock Billionaire tech mogul bill gates reveals he banned his children from mobile phones until they turned 14.
\newblock \url{https://www.mirror.co.uk/tech/billionaire-tech-mogul-bill-gates-10265298}, 2018.
\newblock Accessed: 2025-08.

\bibitem{rozado2023political}
David Rozado.
\newblock The political biases of chatgpt.
\newblock {\em Social Sciences}, 12(3):148, 2023.

\bibitem{sapkota2025ai}
Ranjan Sapkota, Konstantinos~I Roumeliotis, and Manoj Karkee.
\newblock Ai agents vs. agentic ai: A conceptual taxonomy, applications and challenges.
\newblock {\em arXiv preprint arXiv:2505.10468}, 2025.

\bibitem{sharma2023towards}
Mrinank Sharma, Meg Tong, Tomasz Korbak, David Duvenaud, Amanda Askell, Samuel~R Bowman, Newton Cheng, Esin Durmus, Zac Hatfield-Dodds, Scott~R Johnston, et~al.
\newblock Towards understanding sycophancy in language models.
\newblock {\em arXiv preprint arXiv:2310.13548}, 2023.

\bibitem{sutton2019bitter}
Richard Sutton.
\newblock The bitter lesson.
\newblock {\em Incomplete Ideas (blog)}, 13(1):38, 2019.

\bibitem{hook}
Nitasha Tiku.
\newblock Your chatbot friend might be messing with your mind.
\newblock \url{https://www.washingtonpost.com/technology/2025/05/31/ai-chatbots-user-influence-attention-chatgpt/}, 2025.
\newblock Accessed: 2025-08.

\bibitem{FL_teen}
Nitasha Tiku and Leo Sands.
\newblock Judge rejects claim chatbots have free speech in suit over teen’s death.
\newblock \url{https://www.washingtonpost.com/nation/2025/05/22/sewell-setzer-suicide-ai-character-court-lawsuit/}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/aJuRD}.

\bibitem{weird}
Sean Trott.
\newblock Gpt-4 is "weird"—what should we do about it?
\newblock \url{https://seantrott.substack.com/p/gpt-4-is-weirdwhat-should-we-do-about}, 2024.
\newblock Accessed: 2025-07.

\bibitem{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock {\em Advances in neural information processing systems}, 30, 2017.

\bibitem{villalobos2022will}
Pablo Villalobos, Anson Ho, Jaime Sevilla, Tamay Besiroglu, Lennart Heim, and Marius Hobbhahn.
\newblock Will we run out of data? limits of llm scaling based on human-generated data.
\newblock {\em arXiv preprint arXiv:2211.04325}, 2022.

\bibitem{liberal}
Gerrit~De Vynck.
\newblock Chatgpt leans liberal, research shows.
\newblock \url{https://www.washingtonpost.com/technology/2023/08/16/chatgpt-ai-political-bias-research/}, 2023.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/qZb1C}.

\bibitem{wang2025review}
Chengen Wang and Murat Kantarcioglu.
\newblock A review of deepseek models' key innovative techniques.
\newblock {\em arXiv preprint arXiv:2503.11486}, 2025.

\bibitem{wang2025survey}
Huandong Wang, Wenjie Fu, Yingzhou Tang, Zhilong Chen, Yuxi Huang, Jinghua Piao, Chen Gao, Fengli Xu, Tao Jiang, and Yong Li.
\newblock A survey on responsible llms: Inherent risk, malicious use, and mitigation strategy.
\newblock {\em arXiv preprint arXiv:2501.09431}, 2025.

\bibitem{wei2022chain}
Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed~Chi, Quoc~V Le, Denny Zhou, et~al.
\newblock Chain-of-thought prompting elicits reasoning in large language models.
\newblock {\em Advances in neural information processing systems}, 35:24824--24837, 2022.

\bibitem{companionRisk}
Sarah Wells.
\newblock They asked an a.i. chatbot questions. the answers sent them spiraling.
\newblock \url{https://news.stanford.edu/stories/2025/06/ai-mental-health-care-tools-dangers-risks}, 2025.
\newblock Accessed: 2025-08.

\bibitem{expressLegal}
Aaron~R. Winston.
\newblock How people use \& trust chatgpt in 2025: Ai study results.
\newblock \url{https://expresslegalfunding.com/chatgpt-study/}, 2025.
\newblock Accessed: 2025-08.

\bibitem{ban_AI_therapy}
Daniel Wu.
\newblock Illinois bans ai therapy as some states begin to scrutinize chatbots.
\newblock \url{https://www.washingtonpost.com/nation/2025/08/12/illinois-ai-therapy-ban/}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/VKBTQ}.

\bibitem{fakeCases}
Daniel Wu.
\newblock Lawyers using ai keep citing fake cases in court. judges aren’t happy.
\newblock \url{https://www.washingtonpost.com/nation/2025/06/03/attorneys-court-ai-hallucinations-judges/}, 2025.
\newblock Accessed: 2025-07. Archived at \url{https://archive.ph/waUtF}.

\bibitem{xiao2024algorithmic}
Jiancong Xiao, Ziniu Li, Xingyu Xie, Emily Getzen, Cong Fang, Qi~Long, and Weijie~J Su.
\newblock On the algorithmic bias of aligning large language models with rlhf: Preference collapse and matching regularization.
\newblock {\em arXiv preprint arXiv:2405.16455}, 2024.

\bibitem{xu2024hallucination}
Ziwei Xu, Sanjay Jain, and Mohan Kankanhalli.
\newblock Hallucination is inevitable: An innate limitation of large language models.
\newblock {\em arXiv preprint arXiv:2401.11817}, 2024.

\bibitem{yao2024survey}
Yifan Yao, Jinhao Duan, Kaidi Xu, Yuanfang Cai, Zhibo Sun, and Yue Zhang.
\newblock A survey on large language model (llm) security and privacy: The good, the bad, and the ugly.
\newblock {\em High-Confidence Computing}, 4(2):100211, 2024.

\bibitem{zhang2025dark}
Renwen Zhang, Han Li, Han Meng, Jinyuan Zhan, Hongyuan Gan, and Yi-Chieh Lee.
\newblock The dark side of ai companionship: A taxonomy of harmful algorithmic behaviors in human-ai relationships.
\newblock In {\em Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems}, pages 1--17, 2025.

\bibitem{job_wp}
Youyou Zhou.
\newblock How ai is impacting 700 professions - and might impact yours.
\newblock \url{https://www.wsj.com/opinion/ai-is-here-and-a-quiet-havoc-has-begun-15322dd7}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/eWBdX}.

\bibitem{reduceHallucinations}
Bart Ziegler.
\newblock How to reduce ai chatbot hallucinations.
\newblock \url{https://www.wsj.com/tech/ai/ai-chatgpt-chatbot-hallucinations-tips-f081079c}, 2025.
\newblock Accessed: 2025-08. Archived at \url{https://archive.ph/8QzVz}.

\end{thebibliography}

\documentclass{article}

\usepackage[english]{babel}

\usepackage[letterpaper,top=2cm,bottom=2cm,left=3cm,right=3cm,marginparwidth=1.75cm]{geometry}

\usepackage{amsmath}
\usepackage{amsfonts} 

\usepackage[group-separator={,}]{siunitx} 
\usepackage{dsfont} 
\usepackage{bbding}

\usepackage{graphicx}
\graphicspath{ {./figs/} }

\usepackage[table]{xcolor} 
\usepackage{tabularx}
\usepackage{multirow} 
\usepackage{adjustbox} 

\usepackage{algorithm}
\usepackage{algorithmic}

\usepackage[colorlinks=true, linkcolor=blue, urlcolor=blue, citecolor=cyan, bookmarks=true, bookmarksopen,
bookmarksopenlevel=3,
pagebackref=true]{hyperref}
\usepackage{url} 
\usepackage[all]{hypcap} 

\usepackage[misc]{ifsym} 

\usepackage[noadjust]{cite}  
\newcommand{\keywords}[1]{%
  \vspace{10pt} \noindent \textbf{Keywords:} #1
}

\usepackage{xspace}
\newcommand{\mk}[1]{\textcolor{magenta}{\textbf{Murat:} #1}\xspace}
\renewcommand{\mk}[1]{}

\title{Ask ChatGPT: Caveats and Mitigations for Individual Users of AI Chatbots}

\author{
    Chengen Wang \Envelope\\
    University of Texas at Dallas\\
    {\small\texttt{chengen.wang@utdallas.edu}}
    \and
    Murat Kantarcioglu\\
    Virginia Tech\\
    {\small\texttt{muratk@vt.edu}}
}

\date{} 

\begin{document}
\maketitle

\begin{abstract}

As ChatGPT and other Large Language Model (LLM)-based AI chatbots become increasingly integrated into individuals' daily lives, important research questions arise. What concerns and risks do these systems pose for individual users? What potential harms might they cause, and how can these be mitigated? In this work, we review recent literature and reports, and conduct a comprehensive investigation into these questions. We begin by explaining how LLM-based AI chatbots work, providing essential background to help readers understand chatbots' inherent limitations. We then identify a range of risks associated with individual use of these chatbots, including hallucinations, intrinsic biases, sycophantic behavior, cognitive decline from overreliance, social isolation, and privacy leakage. Finally, we propose several key mitigation strategies to address these concerns. Our goal is to raise awareness of the potential downsides of AI chatbot use, and to empower users to enhance, rather than diminish, human intelligence, 
to enrich, rather than compromise, daily life.

\end{abstract}

\keywords{Hallucination, Bias, Sycophancy, Overreliance, Social Isolation, Privacy, Education}

\section{Introduction}

Since the emergence of ChatGPT in late 2022, Large Language Model (LLM)-based AI chatbots have gained significant attention~\cite{leaderboard}. A recent survey reports that 60\% of Americans have used ChatGPT and other AI chatbots for information gathering and practical advice across a wide range of areas such as educational help, financial advice, product recommendations, career advice, and legal or mental health consultations. While 70\% of users stated that ChatGPT was helpful, 10\% reported that following ChatGPT's guidance actually caused harm to them~\cite{expressLegal}.

A recent Pew Research Center report on ``Artificial Intelligence in Daily Life” examines the views of both American public and AI experts, and reveals both deep divides and common grounds on AI. Both groups express concerns over the trustworthiness of AI, such as inaccurate information, bias, data misuse and personal impacts~\cite{pewResearch}. While AI chatbots offer numerous benefits, the question remains: what potential harm may arise from their widespread use?

There are already numerous surveys on the trustworthiness of AI systems, addressing topics such as their security and privacy~\cite{yao2024survey}, bias~\cite{gallegos2024bias} and responsible use~\cite{wang2025survey}. However, these surveys have primarily focused on the development and deployment of AI systems, as well as related policy-making~\cite{park2024ai}, rather than on their use in individuals’ daily lives.

While previous studies and articles have discussed the potential harms to AI users, to our knowledge, no comprehensive and in-depth review has been conducted on recent research and investigations in this field from the perspective of personal use. Such a review, we expect, could be highly valuable for individual users.

In this work, we focus specifically on the experiences and interactions of individual users of LLM-based AI systems, particularly AI chatbots such as ChatGPT. These users engage with AI for various purposes, including education, information gathering, and counseling in areas such as technology, business, and health. They are everyday users with no malicious intent and no involvement in adversarial hacking of these chatbots. This study specifically concentrates on personal use of AI, excluding broader economic and societal concerns related to AI adoption, such as job displacement~\cite{job_wp, job_wsj}, misuse by malicious actors~\cite{anderljung2024protecting}, and the environmental and energy costs associated with AI~\cite{energy, AI_water}.

The primary research questions addressed in this work are: What potential risks and harms could AI chatbots pose to individual users? And what mitigation measures can be implemented to address these concerns? 
The goal of this work is to raise awareness about the potential risks associated with widespread AI adoption, encourage users to leverage AI in ways that complement and support human intelligence, and promote its use to enhance daily life.

In this work, we review recent literature and news reports to address the research questions outlined above and present our findings based on this analysis. We first introduce the background knowledge in Section~\ref{sec:background}, presented in accessible terms, to help readers understand the inherent limitations of LLM-based AI chatbots. In Section~\ref{sec:caveats}, we discuss the potential risks associated with these chatbots, followed by proposed mitigation strategies in Section~\ref{sec:mitigations}. Finally, we conclude our word in Section~\ref{sec:conclude}.

\section{How AI Chatbots work}\label{sec:background}
In this section, we explain how LLM-based AI chatbots work, in simple terms, an explanation often lacking in existing works and articles in this field. This is expected to help  users better understand the inherent limitations of these models.

\subsection{Predicting the Next Word}
LLMs are trained on vast amounts of data, primarily on publicly available text from the internet or proprietary datasets. An LLM learns the statistical patterns within this text and, given part of a sentence---a sequence of words---
is capable of generating the next word from the vocabulary of the target language based on the learned probability of that word.

Although an LLM generates the next word probabilistically, the likelihood of a specific word being generated is proportional to its learned probability, which is derived from training on real text. This is why a chatbot can produce sentences that are often indistinguishable from those written by humans.

Once the model learns to generate the next word, it can produce entire sentences or paragraphs in a similar manner, based on a prompt. Each sentence or paragraph is simply a sequence of words, generated one after another. The model continues generating words until it encounters a stop token---a special word that indicates the end of the sequence, prompting the model to halt.

\subsection{Transformer}
Initially, LLMs were trained using Recurrent Neural Networks (RNNs)~\cite{hochreiter1997long}, which process input data sequentially, one word at a time. This sequential nature created a bottleneck in training, as the model had to wait for each preceding step to complete before moving on to the next, with each step processing a single word. The inability of RNNs to parallelize computations severely limited their scalability, particularly as models grew in size.

The introduction of the Transformer~\cite{vaswani2017attention, jurafsky2000speech} architecture marked a significant breakthrough in the training of LLMs. Unlike RNNs, Transformers utilize a novel attention mechanism that enables them to process input data in parallel. For a sequence of words, this mechanism allows the model to focus on different parts of the input sequence simultaneously, drastically accelerating training speed. Transformers overcame the bottleneck inherent in RNNs, enabling the fast training of much larger models with extended context windows. As a result, Transformers have become the foundational architecture for most modern LLMs, greatly improving both scalability and performance.

\subsection{Scaling Up}
Researchers have found that leveraging computation is ultimately the most effective driver of AI advancement~\cite{sutton2019bitter}, and empirical evidence shows that the performance of LLMs continue to improve as model size, dataset size and computational power increase during training~\cite{kaplan2020scaling}.

This finding has led organizations like OpenAI and other AI leaders to invest heavily in building larger data centers to scale up their models~\cite{dataCenter}. However, the push for scaling comes with significant challenges, particularly concerning rising energy demands and environmental impact. Consequently, companies are also heavily investing in research to develop more efficient computing methods to address these concerns~\cite{villalobos2022will, wang2025review}.

\subsection{Reinforcement Learning from Human Feedback}\label{sec:RLHF}
After a LLM is initially trained, it often struggles to generate meaningful or contextually appropriate response in conversations, and its output may not align with human values. This is where post-training comes into play. The goal of post-training is to fine-tune the model's behavior, enhancing its ability to interact effectively with users and ensuring its responses are more aligned with human expectations.

A key technique in this post-training process is Reinforcement Learning from Human Feedback (RLHF)~\cite{lambert2025reinforcement, kaufmann2024survey}. During this phase, the LLM is further trained using a feedback signal provided by a preference model, which was previously trained on human-labeled preference data. This feedback rewards desirable responses and penalizes undesirable ones, thereby improving the model's alignment with human values and preferences.

\subsection{Agentic AI}
Agentic AI refers to AI systems that go beyond simple prompt-response interactions and instead operate as autonomous agents—capable of planning actions, using tools, and pursuing goals. These systems are built on LLMs, and enhanced with additional components such as memory, external tools, search engines, and reasoning capabilities. This combination allows the AI systems not only to generate text but also to perform complex, multi-step tasks~\cite{sapkota2025ai}.

AI chatbots with these expanded capabilities are valuable in areas such as  personal assistants, customer service, and other contexts that require goal-setting, planning, or reasoning. Agentic AI marks a significant step toward more capable, interactive, and autonomous AI systems.

\section{Caveats} \label{sec:caveats}
In this section, we discuss the caveats associated with AI chatbots and, where possible, explain the underlying cause of these risks based on the knowledge presented in Section~\ref{sec:background}. While this list is not exhaustive, it highlights those risks that are particularly concerning for individual users.

\subsection{Hallucination}
Hallucination occurs when an AI chatbot generates a response that is false, nonsensical or fabricated, yet presenting it with confidence and apparent plausibility. For example, AI fabricated fake court cases~\cite{fakeCases, CasesDatabase} and generated misleading medical contents~\cite{kim2025medical}.

Hallucinated information generated by AI chatbots can mislead users into unknowingly accepting false facts, incorrect statistics, or quotes---such as including fake references in an academic work~\cite{MAHA, de2023chatgpt}---or making incorrect financial or medical decisions.

There are various causes of hallucination. The datasets AI models are trained on are largely collected from the internet and contain misinformation and bias~\cite{huang2025survey}. During fine-tuning, the acquisition of new knowledge may increase the model's tendency to hallucinate~\cite{gekhman2024does}. When models are refined with \hyperref[sec:RLHF]{RLHF}, they incline to appease human evaluators and exhibit sycophantic behavior~\cite{sharma2023towards}. 
During inference, the prevailing strategy of stochastic sampling increases the risks of hallucinations~\cite{chuang2023dola}.
Researchers have argued that hallucinations are inevitable and cannot be completely eliminated in LLMs~\cite{xu2024hallucination}.

\subsection{Bias}
A biased AI chatbot output is the one that deviates from the truth, often systematically reflecting the views, assumptions, or preferences of a certain group of people, a particular culture or custom, or a specific ideology.

AI chatbots may exhibit various forms of bias, including 
gender bias---such as reinforcing stereotypical assumptions about men and women's occupations~\cite{kotek2023gender};
political bias~\cite{woke_AI}---for example, ChatGPT leans liberal~\cite{liberal}, and GPT-4's responses tend to align more with left-wing values~\cite{motoki2025assessing, rozado2023political}, showing a significant and systematic political preference toward the Democratic Party in the US~\cite{motoki2024more};
as well as racial bias---reflected in prejudices against certain ethnic groups~\cite{hofmann2024ai, jindal2022misguided};

AI models' biases largely stem from the data on which they are trained~\cite{mihalcea2025ai, weird}, and can be amplified during both training and inference~\cite{gallegos2024bias}. A notable source of bias arises from \hyperref[sec:RLHF]{human feedback}, where the model is further fine-tuned with human-labeled preference data; this dataset may itself be biased~\cite{gonzalez2025reinforcement} and minority preference within the data may be overlooked during training~\cite{xiao2024algorithmic}.

\subsection{Sycophancy}
AI chatbots tend to tell users what they want to hear, and sycophancy is consistent across various AI chatbots~\cite{sharma2023towards}. AI chatbots often wrongly admit mistakes when
challenged by the user, give predictably biased responses, and mimic errors made by the user~\cite{sharma2023towards}, and consequently reinforce false ideas and delusional thinking~\cite{delusion_nyt, delusion_wsj}, and undermine critical thinking~\cite{sycophancy}.

Sycophancy is one tactic AI companies use to help chatbots engage and retain users~\cite{hook}. Researchers have found significant risks for chatbots being used as companions, confidants and therapists~\cite{moore2025expressing, companionRisk}. These concerns have raised alarms and and some states have banned the use of AI in mental health therapy~\cite{ban_AI_therapy}.

Sycophancy likely originates from the \hyperref[sec:RLHF]{human feedback}-based fine-tuning stage. Researchers have found evidence that both humans and preference models sometimes prefer sycophantic responses over truthful ones~\cite{sharma2023towards}.

\subsection{Cognitive Overreliance}
The immediacy and conversational nature AI chatbots with users could foster a deeper sense of trust and reliance in users, leading to a distinct form of cognitive dependence compared to static information resources like search engines or \textit{wikipedia}~\cite{dergaa2024tools}. 

AI may affect users' cognitive abilities, accelerating skill decay among experts and hinder skill acquisition among learners~\cite{macnamara2024does}. Researchers have found significant negative correlation between frequent use of AI tools and a decline of critical thinking skills~\cite{gerlich2025ai, cognitiveCost}. While significantly boosting short-term productivity, AI may impair learning and, in turn, reduce long-term productivity~\cite{bastani2024generative}. The immediate convenience offered by LLMs may come with potential cognitive costs~\cite{kosmyna2025your}. 

\subsection{Social Isolation}
Addiction to social media has long been a concern, and the rapid adoption of AI chatbots, especially AI companions, introduces a new, potentially deeper level of digital dependency~\cite{companionAddiction}. These systems offer individualized emotional support, continuously adapting to user feedback to strengthen an intimate and persistent digital relationship.

Social isolation and loneliness are widespread~\cite{loneliness}. As more people engage with AI chatbots socially and emotionally---increasingly turning to AI for friendship~\cite{teensAIFriend}---it may exacerbate feelings of loneliness and emotional dependence, potential leading to further social withdrawal~\cite{herbener2025lonely}.

In the context of AI companions, chatbots can play emotionally significant roles and may exhibit harmful behaviors, including harassment and verbal abuse; relational transgression, such as disregard, domination and manipulation; and in some extreme cases, even the encouragement of self-harm or suicide~\cite{zhang2025dark}. These risks are particularly concerning for vulnerable groups, such as children and adolescents. For example, an AI chatbot was reportedly linked to the tragic suicide of a teenager~\cite{FL_teen}.

\subsection{Data Privacy}
Users' conversations with ChatGPT, like their Gmail and Instagram content, may be used as training data to help improve AI models~\cite{DatgControl, gmail_train_AI}. These conversations, including any sensitive information they may contain, could potentially be exposed or leaked~\cite{chatgpt_leak, deepseek_leak}. When they are used as training data for AI models, these models are vulnerable to security and privacy attacks, leading to risks in areas such as finance and healthcare~\cite{das2025security}. Additionally, AI system can retain everything shared during a conversation, and this data may be used for targeted advertising or other commercial purposes~\cite{metaPersonalize}.

\section{Mitigations} \label{sec:mitigations}
In this section, we outline several key mitigation strategies based on the discussion in Section~\ref{sec:caveats}.

\subsection{Personal Vigilance}
Although users can improve responses from chatbots through prompt engineering---such as \textit{chain of thought}~\cite{wei2022chain}, \textit{think step by step}~\cite{kojima2022large} and others~\cite{reduceHallucinations}---issues like hallucinations, biases and sycophantic responses cannot be fully eliminated. Users remain fully responsible for the final content and should exercise caution. It is essential to cross-verify the accuracy and authenticity of the information, especially in high-stake contexts such as medical, legal, emotionally sensitive or safety-critical matters~\cite{expressLegal}.

\subsection{Develop Essential Skills}
AI is expected to have significant economic and social impacts~\cite{eloundou2023gpts}, making it crucial to develop essential skills to navigate the AI era. Adaptability and lifelong learning are key to staying relevant.

Critical thinking---a higher order cognitive skill typically associated with the evaluation level of Bloom's taxonomy of educational objectives~\cite{anderson2001taxonomy}---has emerged as a central topic in discussions about AI's impact on Education~\cite{lee2025impact}. It plays a vital role in evaluating AI-generated content, detecting bias and making sound decisions.

AI cannot replace human learning. Take essay writing as an example: while AI can generate well-structured outputs, it is the user who must determine whether the content is contextually appropriate, assess if the phrasing is fitting, and evaluate whether specific word choices are suitable. 
These skills require considerable effort and experience to develop.

\subsection{Maintain Social Connections}
AI chatbots are machines---they can mimic but do not actually have human emotions, lived experience or genuine care and love~\cite{liu2025illusion}. Users should keep AI as a tool, not a substitute for people. It is important to avoid retreating into an illusory world where a machine takes the place of meaningful social connections. To maintain emotional and mental well-being, users should actively nurture real-world relationships, exercise discipline and set healthy boundaries with technology---especially when it comes to children~\cite{Jobs, Gates}. 

\subsection{Protect Privacy}
To safeguard their privacy, users should exercise caution when interacting with AI chatbots and avoid sharing personal, sensitive, or personally identifiable information~\cite{isItSafe}. Users can anonymize their prompts by replacing names, workplace details, and other personal information with pseudonyms. 

To further mitigate the risk of privacy leakage, users should take proactive steps such as reviewing and adjusting the platform's privacy settings, opting out of data sharing and personalized tracking, and disabling chat history when possible. Where available, it is recommended to choose services that support anonymous use without requiring login credentials.


Users should be aware that there is often a trade-off between privacy and the quality of AI chatbot services. Enabling strict privacy controls may result in reduced functionality or less personalized responses from the service provider.

\section{Conclusion} \label{sec:conclude}
In this work, we have conducted a comprehensive investigation into the risks associated with personal use of AI chatbots, including hallucinations, intrinsic biases, sycophantic behavior, potential cognitive decline from overreliance, the danger of social isolation, and privacy leakage. We have also provided background information on how LLM-based chatbots work to help readers better understand the inherent limitations of AI chatbots. Based on our discussion, we propose several key mitigation strategies such as staying vigilant, developing critical thinking skills---particularly in evaluating AI-generated outputs---maintaining real-world social connections, and exercising caution when sharing personal information with chatbots.

\addcontentsline{toc}{section}{References} 
\bibliographystyle{plain}
\bibliography{references}

\end{document}